# GPU-accelerated LISA parameter estimation with full time domain response

### v2

@@ III.2 Comparison against Fisher analysis
In Fig. 4, we show results for the aligned spin scenario, observing good agreement between the two methods. For the mass parameters, however, the Bayesian posteriors appear slightly narrower than those from the Fisher analysis, potentially due to an artifact of the Fisher initialization. This effect might arise if the sampler locates the peak too quickly, resulting in incomplete sampling of the posterior ellipse. In our other analyses, such as those focused solely on the 22 mode or employing only the AE channels, the Bayesian posteriors appear wider than those obtained from Fisher analysis, as expected. Additionally, it is important to note that the Fisher matrix formalism is particularly sensitive to numerical inaccuracies, which can lead to scenarios where the inverse covariance matrix is not positive semidefinite. Although we did not encounter this issue in the examples presented here, we should be aware that the Fisher posteriors could be affected by numerical errors.
@@
In the precessing case illustrated in Fig. 5, we observe that the posteriors exhibit notable differences, with the Bayesian posteriors consistently wider than those derived from Fisher analysis, which is anticipated. This widening may be attributed to the increased dimensionality associated with precessing systems, requiring the sampler to explore a larger region of the parameter space before locating the correct peak. We should remind that, despite employing Fisher initialization, the priors are wide and allow the sampler to investigate regions outside the Fisher ellipse, potentially influencing the width of the resulting posteriors.
III.3 Comparison equal vs unequal arms
@@
We investigate the impact of incorporating realistic orbits with unequal arms for the LISA constellation, as provided by ESA Martens and Joffre (2021), compared to scenarios assuming equal arms. The equal-arm cases correspond to those discussed in Sec. III.1. For the unequal-arm configuration, we adopt new orbits while keeping the PSDs unchanged. The use of more realistic PSDs will be postponed for future work.
@@
Figs. 6, 7 show the results for the aligned-spin and precessing cases respectively, both including subdominant harmonics. It can be seen that the unequal-arm configuration leads to improved recovery of the ecliptic latitude and distance.
[…]
@@
The work of Katz et al. (2022) also compared the effect of unequal arms with Bayesian inference but on galactic binaries. This study is the first to examine the effect on MBHB with generic spins, representing a significant advancement toward more realistic analyses with LISA. Future work will focus on more systematic investigations of unequal arms, as this area remains largely unexplored.
@@
Figure 6: Posterior distribution comparison between equal and unequal arm orbits for IMRPhenomTHM with 3 TDI channels.
@@
Figure 7: Posterior distribution comparison between equal and unequal arm orbits for IMRPhenomTPHM with 3 TDI channels.
@@ III.4 Comparison of TDI channels
In this section, we present the first analysis of how different TDI channel combinations affect full Bayesian inference in LISA observations. Figs. 8 and 9 display the results for aligned spin and precessing cases, respectively. For each case, injections and recoveries were performed using one (A), two (AE) and three (AET) TDI channels. Channels A and E are constructed to capture the majority of the gravitational wave signal, while the T channel predominantly captures instrumental noise. Excluding the E channel results in broader, less confident posteriors, as it disregards substantial signal information. In contrast, excluding the T channel has little effect under zero-noise injections, though this channel is likely to play a larger role in analyses that include realistic noise, a topic that will be explored in future studies.
@@ III.5 Comparison of TDI generations
We present a novel comparison of parameter estimation posteriors between the first and second TDI generations. To our knowledge, only Wang (2024) has examined this impact but on non-spinning signals. Our work opens the opportunity to explicitly analyze the effect of TDI generations on inference accuracy and assess the potential value of advancing TDI development. We perform the injections and recoveries using three TDI channels (AET) computed for both the first (1.5) and second TDI generations, in aligned spin and precessing cases. We do not display the plots for this case since there are no significant differences between the two TDI generations. For the aligned-spin injection, the posterior distributions overlap almost completely, while for the precessing case, there is a minor difference in the azimuthal angle of the secondary spin. This discrepancy likely reflects the inherent challenges for the sampler in accurately measuring this parameter, as discussed in Sec. III.1, rather than an effect of the TDI generations themselves. Notably, in the precessing case, although the posterior distributions remain similar, the sampler required a significantly greater computational time ($\sim$ 9h) and collected fewer samples overall (see Tables 2 and 3). The increased computational demand associated with the second TDI generation might be due to the more complex structure of its signal response.
@@
Table 2: Median values and error bars for each parameter, with the relative error between the median and the injected value indicated in brackets. The results are compared across runs with and without subdominant harmonics, with and without precessing spins, and with the inclusion of unequal arms (UA).
@@
Table 3: Median values and error bars for each parameter, with the relative error between the median and the injected value indicated in brackets. The results are compared across runs with channels A and AE and the second TDI generation for AET.
@@ IV Conclusions
We have performed the first full Bayesian parameter estimation study for LISA MBHBs employing a time domain response and including the effects of subdominant harmonics and precession. The time domain approach offers a more robust theoretical formalism compared to the approximated transfer functions formalism in the Fourier domain. It facilitates the straightforward inclusion of various physical effects in the waveform, such as precession, eccentricity, and memory, and provides a basis to test and validate the accuracy of the Fourier domain formalism.
Until now, time domain methods have been computationally prohibitive for LISA parameter estimation. We have significantly reduced the computational cost of waveform generation and likelihood evaluation to make this approach feasible. This was achieved using a new Python implementation of the IMRPhenomT waveform family with GPU support, which parallelizes the computation of each waveform’s time array. Additionally, the precession module in this model now applies a more efficient polarization calculation method, distinct from that used by the LVK collaboration. The LISA arm response and TDI channels are computed directly in the time domain through a GPU-accelerated version of the LISA Data Challenge tools. Both Fourier transforms and inner products required for likelihood calculations are processed on the GPU, with the final likelihood value passed to the Bayesian inference package Bilby for stochastic sampling and likelihood maximization across the parameter space.
We utilized the ptemcee sampler with Fisher initialization and broad uniform priors to conduct zero-noise injections for both aligned-spin and precessing systems. Our analysis investigates the influence of subdominant harmonics, as well as different TDI channels and TDI generations, on parameter recovery. To further validate these findings, we compared the resulting Bayesian posteriors with those derived from the standard Fisher analysis, observing overall consistency with the expected theoretical behaviour across all parameters.
@@
This work represents the first Bayesian inference study to explicitly assess the impact of unequal arms, different TDI channels and TDI generations on LISA parameter estimation for MBHBs. Results indicate that the realistic orbits with unequal arms improve the measurement of the sky localization and distance. We also observe that the T channel does not influence posterior distributions, which aligns with expectations, as this channel is designed to be noise-dominated, and our injected signals are noise-free. Similarly, no significant differences emerged between TDI generations in the current setup. However, this does not rule out the need for further development of TDI channels and generations, since the influence of TDI configurations is likely to be much more pronounced in analyses involving realistic noise. Our next objective is to extend this study by incorporating gaussian noise into the signal, which remains unexplored even in Fourier domain approaches.
Our analysis was executed on a single GPU, but future works will involve running multiple Markov chains Monte Carlo in parallel across several GPUs, which will enhance computational efficiency. Additionally, we aim to introduce GPU support for alternative precessing prescriptions within the IMRPhenomT model, potentially improving the accuracy of certain parameter measurements. Current development also includes incorporating eccentricity and the memory effect into our model, which our parameter estimation pipeline can accommodate seamlessly. This enhancement would enable the first studies of these effects on LISA parameter estimation, broadening our capacity to gain richer physical information from the signal and advancing our understanding of the underlying astrophysical processes.
We recognize that optimizing sampler settings could further reduce sampling time and improve posterior convergence; however, this will be addressed in future studies, alongside trials of alternative samplers to evaluate their effectiveness.
One current limitation is the necessity to initialize the sampler with Fisher estimates to enhance convergence speed. This requirement could potentially be addressed by further reducing the computational cost associated with likelihood evaluations. In this regard, we are exploring more efficient algorithms for waveform and likelihood computation, including non-uniform time grids and interpolation methods such as Multibanding, Relative Binning or FANTA [^6] FANTA (Fourier Analytical Transform of a piecewise Approximant) leverages the Fourier transform’s linearity to analytically integrate a simplified form of the waveform function across each time interval.. By implementing these acceleration techniques, we also aim to extend this framework to lower-mass systems with longer-duration signals, such as those from stellar-mass black holes.
The next steps and challenges can be summarized in
@@
  • Optimization of sampler settings for general initialization
@@
  • Further speed-up of waveform and likelihood evaluation with non-uniform grids algorithms.
@@ Acknowledgements
We deeply thank Stanislav Babak for his support and encouragement throughout this project. We thank Jorge Valencia for finding a bug in the LISA response code and for useful discussions. CG is supported by the Swiss National Science Foundation (SNSF) Ambizione grant PZ00P2_223711 and was also supported by the Sinergia grant 213497. ST is supported by the Swiss National Science Foundation Ambizione Grant Number : PZ00P2-202204. We gratefully acknowledge support from the CNRS/IN2P3 Computing Center (Lyon - France) for providing computing and data-processing resources needed for this work.

### v3

@@ III.2 Comparison against Fisher analysis
In Fig. 5, we show results for the aligned spin scenario, observing good agreement between the two methods. For the mass parameters, however, the Bayesian posteriors appear slightly narrower than those from the Fisher analysis, potentially due to an artifact of the Fisher initialization. This effect might arise if the sampler locates the peak too quickly, resulting in incomplete sampling of the posterior ellipse. In our other analyses, such as those focused solely on the 22 mode or employing only the AE channels, the Bayesian posteriors appear wider than those obtained from the Fisher analysis, as expected. Additionally, it is important to note that the Fisher matrix formalism is particularly sensitive to numerical inaccuracies, which can lead to scenarios where the inverse covariance matrix is not positive semidefinite. Although we did not encounter this issue in the examples presented here, we should be aware that the Fisher posteriors could be affected by numerical errors.
@@
In the precessing case illustrated in Fig. 6, we observe that the posteriors exhibit notable differences, with the Bayesian posteriors consistently wider than those derived from the Fisher analysis, which is anticipated. This widening may be attributed to the increased dimensionality associated with precessing systems, requiring the sampler to explore a larger region of the parameter space before locating the correct peak. We should remind that, despite employing Fisher initialization, the priors are wide and allow the sampler to explore regions outside the Fisher ellipse, potentially influencing the width of the resulting posteriors.
III.3 Comparison equal vs unequal-arms
@@
We investigate the impact of incorporating realistic orbits with unequal-arms for the LISA constellation, as provided by ESA Martens and Joffre (2021), compared to scenarios assuming equal-arms. The equal-arm cases correspond to those discussed in Sec. III.1. For the unequal-arm configuration, we adopt new orbits while keeping the PSDs unchanged. The use of more realistic PSDs will be postponed for future work.
@@
Figs. 7, 8 show the results for the aligned-spin and precessing cases respectively, both including subdominant harmonics. It can be seen that the unequal-arm configuration leads to improved recovery of the ecliptic latitude and distance.
Variations in the LISA arm lengths are intrinsically linked to the sky location of the source (see Eq. 8), meaning that additional amplitude and phase time-dependent modulations introduced by LISA’s breathing motion help to resolve degeneracies and improve measurement precision. The improved sky location together with the amplitude modulations helps to better constrain the distance. The angular resolution of LISA depends on sky location Cutler (1998), and changes in ecliptic latitude and longitude influence the signal differently depending on the specific sky location parameters. For the values considered here, a shift in latitude of $1^{\circ}$ causes amplitude and phase variations in the TDI channels an order of magnitude larger than a shift in longitude, highlighting greater sensitivity in the latitude. This role may vary across different sky locations, with the longitude becoming more sensitive. A comprehensive study of this behaviour across the celestial sphere is deferred to future work. Additionally, a more detailed analysis of how variations in orbital trajectories influence the recovery of the parameters is also left for subsequent investigation.
@@
The work of Katz et al. (2022) also compared the effect of unequal-arms with Bayesian inference but on galactic binaries. This study is the first to examine the effect on MBHB with generic spins, representing a significant advancement toward more realistic analyses with LISA. Future work will focus on more systematic investigations of unequal-arms, as this area remains largely unexplored.
@@
Figure 7: Posterior distribution comparison between equal and unequal-arm orbits for IMRPhenomTHM with 3 TDI channels.
@@
Figure 8: Posterior distribution comparison between equal and unequal-arm orbits for IMRPhenomTPHM with 3 TDI channels.
@@ III.4 Comparison of TDI channels
In this section, we present the first analysis of how different TDI channel combinations affect full Bayesian inference in LISA observations. Figs. 9 and 10 display the results for aligned spin and precessing cases, respectively. For each case, injections and recoveries were performed using one (A), two (AE) and three (AET) TDI channels. Channels A and E are constructed to capture the majority of the gravitational wave signal, while the T channel predominantly captures instrumental noise. Excluding the E channel results in broader, less confident posteriors, as it disregards substantial signal information. In contrast, excluding the T channel has little effect under zero-noise injections, though this channel is likely to play a larger role in analyses that include realistic noise, a topic that will be explored in future studies.
@@ III.5 Comparison of TDI generations
We present a novel comparison of parameter estimation posteriors between the first (1.5) and second TDI generations. To our knowledge, only Wang (2024) has examined this impact but on non-spinning signals. Our work opens the opportunity to explicitly analyze the effect of TDI generations on inference accuracy and assess the potential value of advancing TDI development. We perform the injections and recoveries using three TDI channels (AET) computed for both the first (1.5) and second TDI generations, in aligned spin and precessing cases. We do not display the plots for this case since there are no significant differences between the two TDI generations, but the recovered values are provided in Tab. 3. For the aligned-spin injection, the posterior distributions overlap almost completely, while for the precessing case, there is a minor difference in the azimuthal angle of the secondary spin. This discrepancy likely reflects the inherent challenges for the sampler in accurately measuring this parameter, as discussed in Sec. III.1, rather than an effect of the TDI generations themselves. Notably, in the precessing case, although the posterior distributions remain similar, the sampler required a significantly greater computational time ($\sim$ 9h) and collected fewer samples overall (see Tables 2 and 3). The increased computational demand associated with the second TDI generation might be due to the more complex structure of its signal response.
@@
Table 2: Median values and error bars for each parameter, with the relative error between the median and the injected value indicated in brackets. The results are compared across runs with and without subdominant harmonics, with and without precessing spins, and with the inclusion of unequal-arms (UA).
@@
Table 3: Median values and error bars for each parameter, with the relative error between the median and the injected value indicated in brackets. The results are compared across runs with channels A and AE and the second TDI generation (G2) for AET.
@@ IV Conclusions
We have performed the first full Bayesian parameter estimation study for LISA MBHBs employing a time domain response and including the effects of subdominant harmonics and precession. The time domain approach offers a more robust theoretical formalism compared to the approximated transfer functions formalism used in the Fourier domain. It facilitates a straightforward inclusion of various physical effects in the waveform, such as precession, eccentricity or memory, and provides a basis to test and validate the accuracy of the Fourier domain formalism.
Until now, time domain methods have been computationally prohibitive for LISA parameter estimation. We have significantly reduced the computational cost of waveform generation and likelihood evaluation to make this approach feasible. This was achieved using a new Python implementation of the IMRPhenomT waveform family with GPU support, which parallelizes the computation of each waveform’s time array. Additionally, the precession module in this waveform model now applies a more efficient algorithm to compute the polarizations, distinct from that in the LALSuite implementation used by the LVK. The LISA arm response and TDI channels are computed directly in the time domain also through a GPU-accelerated version of the LISA Data Challenge tools. Both the Fourier transforms and the inner product calculations required for the likelihood evaluation are processed on the GPU, with the final likelihood value passed to the Bayesian inference package Bilby for stochastic sampling and likelihood maximization across the parameter space.
We utilized the ptemcee sampler with Fisher initialization and broad uniform priors to conduct zero-noise injections both for aligned-spin and precessing systems. Our analysis investigates the influence of subdominant harmonics and unequal-arm orbits, as well as different TDI channels and TDI generations on parameter recovery. To further validate these findings, we compared the resulting Bayesian posteriors with those derived from the standard Fisher analysis, observing overall consistency with the expected theoretical behaviour across all parameters.
@@
This work represents the first Bayesian inference study to explicitly assess the impact of unequal-arms, different TDI channels and TDI generations on LISA parameter estimation for MBHBs. Results indicate that the realistic orbits with unequal-arms improve the measurement of the sky localization and distance. We also observe that the T channel does not influence posterior distributions, which aligns with expectations, as this channel is designed to be noise-dominated, and our injected signals are noise-free. Similarly, no significant differences emerged between TDI generations in the current setup. However, this does not rule out the need for further development of TDI channels and generations, since the influence of TDI configurations is likely to be much more pronounced in analyses involving realistic noise. Our next objective is to extend this study by incorporating Gaussian noise into the signal, which remains unexplored even in Fourier domain approaches.
Our analysis was executed on a single GPU, but future work will involve running multiple Markov chains Monte Carlo in parallel across several GPUs, which will enhance computational efficiency. Additionally, we aim to introduce GPU support for alternative precessing prescriptions within the IMRPhenomT model, potentially improving the accuracy of certain parameter measurements. Current development also includes incorporating eccentricity and the memory effect into our model, which our parameter estimation pipeline can accommodate seamlessly. This enhancement would enable the first studies of these effects on LISA parameter estimation, broadening our capacity to gain richer physical information from the signal and advancing our understanding of the underlying astrophysical processes.
We recognize that optimizing sampler settings could further reduce sampling time and improve posterior convergence; however, this will be addressed in future studies, alongside trials of alternative samplers to evaluate their effectiveness. One current limitation is the necessity of initializing the sampler with Fisher estimates to enhance convergence speed. This requirement could potentially be addressed by further reducing the computational cost associated with likelihood evaluations. In this regard, we are exploring more efficient algorithms for waveform and likelihood computation, including nonuniform time grids and interpolation methods such as Multibanding, Relative Binning or FANTA [^6] FANTA (Fourier Analytical Transform of a piecewise Approximant) leverages the Fourier transform’s linearity to analytically integrate a simplified form of the waveform function across each time interval.. By implementing these acceleration techniques, we also aim to extend this framework to lower-mass systems with longer-duration signals, such as those from stellar-mass black holes.
The next steps and challenges can be summarized as
@@
  • Optimization of sampler settings for general sampler initialization.
@@
  • Further speed-up of waveform and likelihood evaluation with nonuniform grids algorithms.
Data Availability
Data supporting the findings of this article are openly available García-Quirós (2025).
@@ Acknowledgements
We thank Jorge Valencia for finding a bug in the LISA response code and for useful discussions. CG is supported by the Swiss National Science Foundation (SNSF) Ambizione grant PZ00P2_223711 and was also supported by Sinergia grant 213497. ST is supported by the Swiss National Science Foundation Ambizione Grant Number : PZ00P2-202204. We gratefully acknowledge support from the CNRS/IN2P3 Computing Center (Lyon - France) for providing computing and data-processing resources needed for this work.